# Whitney Approximation: domains and bounds

### v1

@@ Introduction
In this paper we let $a$, $b$, $s$, $t$ range over ${\mathbb{R}}$, $m$, $n$ over =${\mathbb{N}}=\{0,1,2,\dots\}$, $r$ over ${}{{\mathbb{N}}\cup\{\infty\}}$, and $z$ over ${\mathbb{C}}$. Let $I=(\alpha,\beta)$, where $-\infty\leqslant\alpha<\beta\leqslant\infty$, be an open interval. Every $\mathcal{C}^{r}$-function ${I\to{\mathbb{R}}}$ and its derivatives can be arbitrarily well approximated by an analytic function and its derivatives; more precisely:
@@
This theorem is implicit in the proof of a (more general) lemma in Whitney’s seminal 1934 paper on extending multivariate differentiable functions [25]. In [3, Appendix A] we included a self-contained proof of this fact, which makes it apparent that there is an open subset $U=U_{I}$ of ${\mathbb{C}}$ which contains $I$, independent of $f$, such that the function $g$ in the conclusion of the theorem extends to a holomorphic function ${U\to{\mathbb{C}}}$. If $I$ is bounded, then the open set $U$ obtained through the proof of Whitney’s theorem as in loc. cit. is also bounded. (See [3, Corollary A.4].) However, if $I={\mathbb{R}}$, then this procedure yields $U={\mathbb{C}}$, so $g$ can be taken as the restriction of an entire function. Independently of Whitney, this fact had also been discovered by Hoischen [14], and for $r=0$ had already been shown by Carleman [8] in 1927 and for $r=1$ by Kaplan [15] in 1955. (See [20] for the history of such approximation theorems, going back to Weierstrass.) By a theorem of Gauthier and Kienzle [13, Theorem 1.3], in the case ${r<\infty}$ one can always take $U=I\cup({\mathbb{C}}\setminus{\mathbb{R}})$. Our first main result strengthens this fact by showing that it is possible to choose $U={\mathbb{C}}\setminus\{\alpha,\beta\}$ (even when $r=\infty$):
@@
$-1+\cos\beta(t)<\frac{g(t)}{\varepsilon(t)}<1+\cos\beta(t),$
so between any two reals $a$, $b$ with $\cos\beta(a)=1$, $\cos\beta(b)=-1$ there is a zero of $g$. This implies that each interval $[-t,t]$ ($t\geqslant 0$) contains $\geqslant(2/\pi)\beta(t)-2$ zeros of $g$. Using Jensen’s Formula [6, 1.2.1] we obtain a $C\in{\mathbb{R}}$ such that
@@
[…]
@@
Suppose $r<\infty$. Then for each $f\in\mathcal{C}^{r+1}({\mathbb{R}})$ and $\varepsilon\in\mathcal{C}({\mathbb{R}}^{\geqslant})$ such that $\varepsilon>0$ is strictly decreasing and convex with $\varepsilon(t)\to 0$ as $t\to\infty$, there are $C,D\geqslant 1$ and a real entire function $g$ such that $\lvert(f-g)^{(n)}(t)\rvert<\varepsilon(\lvert t\rvert)$ for $n\leqslant r$, and
@@
Let $\varepsilon\in\mathcal{C}^{1}({\mathbb{R}}^{\geqslant})$ be decreasing with $\varepsilon>0$, such that $\lim\limits_{t\to\infty}t\varepsilon^{\prime}(t)/\varepsilon(t)\in{\mathbb{R}}^{\leqslant}$ exists, and let $f\in\mathcal{C}^{1}({\mathbb{R}})$. Then by [2, Theorem 6] there are a $C\in{\mathbb{R}}^{>}$ and a real entire function $g$ such that $\lvert f(t)-g(t)\rvert<\varepsilon(\lvert t\rvert)$ for each $t$ and
@@
If $\lambda_{m}(g)<\infty$ for some $m$, then $g$ is said to have finite index, and in this case the smallest such $m$ is called the index of $g$. Note that if $g$ has finite index $m$, then ${m\geqslant 1}$. By convention, constant functions ${\mathbb{C}}\to{\mathbb{C}}$ have index 00. The entire functions of index $1$ are exactly the non-constant polynomial functions. The entire functions of index $\leqslant 2$ are also known as the entire functions of finite order, and  $\lambda_{2}(g)$ is called the order of $g$. (See [6, Chapter 2].) By [22, Theorem 1], if $g$ has index $m\geqslant 2$, then $\lambda=\lambda_{m}(g)$ may be computed from the Taylor coefficients $g_{n}:=g^{(n)}(0)/n!$ of $g$ at 00 as follows:
@@
Let $f\in\mathcal{C}^{1}({\mathbb{R}})$ and $M\in{\mathbb{R}}^{\geqslant 1}$ be such that $\lVert f\rVert,\lVert f^{\prime}\rVert\leqslant M$, and $\varepsilon>0$. Then by the remark following Corollary 2, there is a $C\in{\mathbb{R}}^{>}$ and a real entire function $g$ such that $\lvert(f-g)(t)\rvert<\varepsilon$ for all $t$ and
@@
[…]
@@
[…]
@@ 2. Proof of Theorem 1
See, e.g., [9, §3.3] for basic facts about the $B_{mn}$. For example, they have coefficients in ${\mathbb{N}}$, with $B_{00}=1$, $B_{0n}=0$ for $n\geqslant 1$, and for $1\leqslant m\leqslant n$, $B_{mn}$ is homogeneous of degree $m$. Let {${}{n}{m}$ denote the Stirling numbers of the second kind, that is, the number of equivalence relations on an $n$-element set with exactly $m$ equivalence classes; cf. [4, p. 576]. They obey the recurrence relations
@@
We now prove Theorem 1. Thus let $I$ be an open interval, $f\in\mathcal{C}^{r}(I)$, and ${\varepsilon,\rho\in\mathcal{C}(I)}$ where $\varepsilon>0$, $r\geqslant\rho\geqslant 0$. We first assume that $I$ is bounded, say $I=(a,b)$ where $a<b$, and we need to show the existence of a holomorphic function $g\colon{\mathbb{C}}\setminus\{a,b\}\to{\mathbb{C}}$ such that $g(I)\subseteq{\mathbb{R}}$ and $\lvert({f-g})^{(n)}(t)\rvert<\varepsilon(t)$ for all $t\in I$ and $n\leqslant\rho(t)$. It is straightforward to arrange that $I=(-1,1)$: Set $\alpha:=\frac{b-a}{2}$, $\beta:=\min\{\alpha,1\}$, and $J:=(-1,1)$, and consider the holomorphic bijection $\phi\colon{\mathbb{C}}\to{\mathbb{C}}$, $\phi(z):=\alpha\cdot z+\textstyle\frac{1}{2}(a+b)$, with compositional inverse $\phi^{\operatorname{inv}}$. Put $f_{0}:=(f\circ\phi)|_{J}\colon J\to{\mathbb{R}}$ and let $\varepsilon_{0},\rho_{0}\in\mathcal{C}(J)$ be given by $\varepsilon_{0}(s):=\beta^{\rho(\phi(s))}\varepsilon(\phi(s))$ and $\rho_{0}(s):=\rho(\phi(s))$, respectively. Suppose we have a holomorphic function $g_{0}\colon{\mathbb{C}}\setminus\{\pm 1\}\to{\mathbb{C}}$ such that $g_{0}(s)\in{\mathbb{R}}$ and
@@
Hence replacing $f$, $\varepsilon$, $\rho$ by $f_{0}$, $\varepsilon_{0}$, $\rho_{0}$, respectively, we may arrange $I=(-1,1)$, which we assume from now on. The rational function
@@
This shows Theorem 1 in the case of a bounded interval. Now suppose $I\subseteq{\mathbb{R}}$ is unbounded, say $I=(a,\infty)$. Replacing $a$, $f$, $\varepsilon$, $\rho$ by 00, $t\mapsto f(t+a)$, $t\mapsto\varepsilon(t+a)$, and $t\mapsto\rho(t+a)$, respectively, we first arrange $a=0$, so $I={\mathbb{R}}^{>}$. Now consider the rational function
@@
Next we prove Corollary 1. Let $I$, $f$, $\varepsilon$, $\rho$ be as in the corollary. The case where $I$ is open is covered by Theorem 1, and the case where $I$ is closed and bounded is a consequence of the Weierstrass Approximation Theorem [18, p. 33]. Suppose ${I=[\alpha,\beta)}$ where $-\infty<\alpha<\beta\leqslant\infty$. (The remaining case where $I=(\alpha,\beta]$ with $-\infty\leqslant\alpha<\beta<\infty$ is treated in a similar way.) Then $\operatorname{fr}(I)=\{\beta\}$. We set $I_{1}:=(-\infty,\beta)$ and now use a well-known fact:
@@ 3. Bounds on the Derivatives of Bump Functions
By a bump function we mean here a $\mathcal{C}^{\infty}$-function $\alpha\colon{\mathbb{R}}\to{\mathbb{R}}$ which is 00 on $(-\infty,0]$, strictly increasing on $[0,1]$, and $1$ on $[1,+\infty)$. Such bump functions play an important role in many constructions in analysis, e.g., partitions of unity. In this section we give a construction of a bump function $\alpha$ with controlled growth of its derivatives $\alpha^{(n)}$. As an auxiliary result we first establish the following formula for the derivatives of reciprocals of nonzero elements of a differential field $K$:
@@
We work in the setting of [4, Chapter 12] and assume familiarity with the concepts and the notations introduced there. Let $\phi$ range over $K^{\times}$ and $i$$j$$k$$l$ over ${\mathbb{N}}$. Put
@@
By the generalized Leibniz Rule, $\phi\mapsto h_{\phi}\colon K^{\times}\to K[[z]]^{\times}$ is a group morphism, so we obtain a group morphism $\phi\mapsto[h_{\phi}]\colon K^{\times}\to\mathfrak{tr}_{K}$ whose image is contained in the Appell group $\mathcal{A}$ over $K$ (cf. [4, p. 565]). For $i\leqslant j$ we have $[h_{\phi}]_{ij}={j\choose i}\phi^{(j-i)}/\phi$. Since $\mathcal{A}\subseteq 1+\mathfrak{tr}_{K}^{1}$, we have $\big{(}[h_{\phi}]-1\big{)}^{k}\in\mathfrak{tr}_{K}^{k}$ for $k\geqslant 1$, and
$[h_{\phi^{-1}}]=[h_{\phi}]^{-1}=\sum_{k}(-1)^{k}\big{(}[h_{\phi}]-1\big{)}^{k}=\sum_{k}(-1)^{k}\left(\sum_{l=0}^{k}{k\choose l}(-1)^{k-l}[h_{\phi^{l}}]\right).$
@@
Let $\alpha$, $c$, $d$ be as in Proposition 3.2, where we may assume $c\geqslant 1$. With $10^{0}:=$ 1, put $C_{n}:=cn^{dn}$, so $1=C_{0}\leqslant C_{1}\leqslant\cdots$. For $a<b$ in ${\mathbb{R}}$, we define the increasing $\mathcal{C}^{\infty}$-function $\alpha_{a,b}\colon{\mathbb{R}}\to{\mathbb{R}}$ by
@@
Then $\alpha_{a,b,a_{*},b_{*}}(t)=0$ if $t\notin[a,b_{*}]$, $\alpha_{a,b,a_{*},b_{*}}(t)=1$ if $t\in[b,a_{*}]$, and $\alpha_{a,b,a_{*},b_{*}}$ is increasing on $[a,b]$ and decreasing on $[a_{*},b_{*}]$. Suppose $\varepsilon,\varepsilon_{*}\leqslant 1$. Then by (3.6),
[…]
In particular, for suitable reals $c_{1},d_{1}\geqslant 1$ (depending only on $b-a$$b_{*}-a_{*}$$c$$d$) we have $\lVert\alpha_{a,b,a_{*},b_{*}}\rVert_{n}\leqslant c_{1}n^{d_{1}n}$ for each $n$.
@@ 4. Weierstrass Transforms
Recall that the support $\operatorname{supp}f$ of a function $f\colon{\mathbb{R}}\to{\mathbb{R}}$ is the closure in ${\mathbb{R}}$ of the set ${\big{\{}t\in U:\,f(t)\neq 0\big{\}}}$. Let $f\in\mathcal{C}^{m}({\mathbb{R}})$ be such that $\operatorname{supp}f$ is bounded; let also $\lambda$ range over ${\mathbb{R}}^{>}$. The function ${f_{\lambda}\colon{\mathbb{R}}\to{\mathbb{R}}}$ given by
@@
is known as the (generalized) Weierstrass transform with parameter $\lambda$; below we sometimes denote $f_{\lambda}$ also by $W_{\lambda}(f)$. We could have replaced the bounds $\infty$- $\infty$ in this integral by any $a$$b$ such that $\operatorname{supp}(f)\subseteq[a,b]$. A change of variables gives

### v2

@@ Introduction
In this paper we let $a$, $b$, $s$, $t$ range over $\mathbb{R}$}, $m$, $n$ over ${\mathbb{N}}=\{0,1,2,\dots\}$, $r$ over $\mathbb{N}$$\cup\{\infty\}$}, and $z$ over $\mathbb{C}$}. Let $I=(\alpha,\beta)$, where $\infty\leqslant\alpha<\beta\leqslant\infty$, be an open interval. Every $\mathcal{C}^{r}$-function $I\to{\mathbb{R}}$} and its derivatives can be arbitrarily well approximated by an analytic function and its derivatives; more precisely:
@@
This theorem is implicit in the proof of a (more general) lemma in Whitney’s seminal 1934 paper on extending multivariate differentiable functions [25]. In [3, Appendix A] we included a self-contained proof of this fact, which makes it apparent that there is an open subset $U=U_{I}$ of $\mathbb{C}$} which contains $I$, independent of $f$, such that the function $g$ in the conclusion of the theorem can be taken as the restriction of a holomorphic function $U\to{\mathbb{C}}$}. If $I$ is bounded, then the open set $U$ obtained through the proof of Whitney’s theorem as in loc. cit. is also bounded. (See [3, Corollary A.4].) However, if $I={\mathbb{R}}$, then this procedure yields $U={\mathbb{C}}$, so $g$ can be taken as the restriction of an entire function. Independently of Whitney, this fact had also been discovered by Hoischen [14], and for $r=0$ had already been shown by Carleman [8] in 1927 and for $r=1$ by Kaplan [15] in 1955. (See [20] for the history of such approximation theorems, going back to Weierstrass.) By a theorem of Gauthier and Kienzle [13, Theorem 1.3], in the case $r<\infty$} one can always take $U=I\cup({\mathbb{C}}\setminus{\mathbb{R}})$. Our first main result strengthens this fact by showing that it is possible to choose $U={\mathbb{C}}\setminus\{\alpha,\beta\}$ (even when $r=\infty$):
@@
$1+\cos\beta(t)<\frac{g(t)}{\varepsilon(t)}<1+\cos\beta(t),$
so between any two reals $a$, $b$ with $\cos\beta(a)=1$, $\cos\beta(b)=-1$ there is a zero of $g$. This implies that each interval [t,t][-t,t] ($t\geqslant 0$) contains $\geqslant(2/\pi)\beta(t)-2$ zeros of $g$. Using Jensen’s Formula [6, 1.2.1] we obtain a $C\in{\mathbb{R}}$ such that
@@
$\lVert f\rVert_{t;\,\rho}:=\sup\big{\{}\lvert f^{(n)}(s)\rvert:\lvert s\rvert\leqslant t,\ n\leqslant\rho\big{\}}\in{\mathbb{R}}^{\geqslant},\qquad\lVert f\rVert_{t}:=\lVert f\rVert_{t;\,0}.$
@@
[…]
@@
[…]
@@
If $\lambda_{m}(g)<\infty$ for some $m$, then $g$ is said to have finite index, and in this case the smallest such $m$ is called the index of $g$. Note that if $g$ has finite index $m$, then $m\geqslant 1$}. By convention, constant functions $\mathbb{C}$$\to{\mathbb{C}}$ have index 0. The entire functions of index $1$ are exactly the non-constant polynomial functions. The entire functions of index $\leqslant 2$ are also known as the entire functions of finite order, and  $\lambda_{2}(g)$ is called the order of $g$. (See [6, Chapter 2].) By [22, Theorem 1], if $g$ has index $m\geqslant 2$, then $\lambda=\lambda_{m}(g)$ may be computed from the Taylor coefficients $g_{n}:=g^{(n)}(0)/n!$ of $g$ at 0 as follows:
@@
Let $f\in\mathcal{C}^{1}({\mathbb{R}})$ and $M\in{\mathbb{R}}^{\geqslant 1}$ be such that $\lVert f\rVert,\lVert f^{\prime}\rVert\leqslant M$, and $\varepsilon\in{\mathbb{R}}^{>}$. Then by the remark following Corollary 2 (applied to the constant function $\varepsilon$), there is a $C\in{\mathbb{R}}^{>}$ and a real entire function $g$ such that $\lvert(f-g)(t)\rvert<\varepsilon$ for all $t$ and
@@
$\lVert f\rVert_{t;\,\rho}:=\sup\big{\{}\lvert f^{(n)}(s)\rvert:0\leqslant s\leqslant t,\ n\leqslant\rho\big{\}}\in{\mathbb{R}}^{\geqslant},\qquad\lVert f\rVert_{t}:=\lVert f\rVert_{t;\,0}.$
@@
(0.1) $\lVert fg\rVert_{S;\,m}\ \leqslant\ 2^{m}\lVert f\rVert_{S;\,m}\lVert g\rVert_{S;\,m}.$
@@ 2. Proof of Theorem 1
[…]
@@
We now prove Theorem 1. Thus let $I$ be an open interval, $f\in\mathcal{C}^{r}(I)$, and $\varepsilon,\rho\in\mathcal{C}(I)$} where $\varepsilon>0$, $r\geqslant\rho\geqslant 0$. We first assume that $I$ is bounded, say $I=(a,b)$ where $a<b$, and we need to show the existence of a holomorphic function $g\colon{\mathbb{C}}\setminus\{a,b\}\to{\mathbb{C}}$ such that $g(I)\subseteq{\mathbb{R}}$ and $\lvert({f-g})^{(n)}(t)\rvert<\varepsilon(t)$ for all $t\in I$ and $n\leqslant\rho(t)$. It is straightforward to arrange that I=(1,1)I=(-1,1): Set $\alpha:=\frac{b-a}{2}$, $\beta:=\min\{\alpha,1\}$, and J:=(1,1)J:=(-1,1), and consider the holomorphic bijection $\phi\colon{\mathbb{C}}\to{\mathbb{C}}$, $\phi(z):=\alpha\cdot z+\textstyle\frac{1}{2}(a+b)$, with compositional inverse $\phi^{\operatorname{inv}}$. Put $f_{0}:=(f\circ\phi)|_{J}\colon J\to{\mathbb{R}}$ and let $\varepsilon_{0},\rho_{0}\in\mathcal{C}(J)$ be given by $\varepsilon_{0}(s):=\beta^{\rho(\phi(s))}\varepsilon(\phi(s))$ and $\rho_{0}(s):=\rho(\phi(s))$, respectively. Suppose we have a holomorphic function $g_{0}\colon{\mathbb{C}}\setminus\{\pm 1\}\to{\mathbb{C}}$ such that $g_{0}(s)\in{\mathbb{R}}$ and
@@
Hence replacing $f$, $\varepsilon$, $\rho$ by $f_{0}$, $\varepsilon_{0}$, $\rho_{0}$, respectively, we may arrange I=(1,1)I=(-1,1), which we assume from now on. The rational function
@@
This shows Theorem 1 in the case of a bounded interval. Now suppose $I\subseteq{\mathbb{R}}$ is unbounded, say $I=(a,\infty)$. Replacing $a$, $f$, $\varepsilon$, $\rho$ by 0, $t\mapsto f(t+a)$, $t\mapsto\varepsilon(t+a)$, and $t\mapsto\rho(t+a)$, respectively, we first arrange $a=0$, so $I={\mathbb{R}}^{>}$. Now consider the rational function
@@
Next we prove Corollary 1. Let $I$, $f$, $\varepsilon$, $\rho$ be as in the corollary. The case where $I$ is open is covered by Theorem 1, and the case where $I$ is closed and bounded is a consequence of the Weierstrass Approximation Theorem [18, p. 33]. Suppose $I=[\alpha,\beta)$} where $\infty<\alpha<\beta\leqslant\infty$. (The remaining case where $I=(\alpha,\beta]$ with $\infty\leqslant\alpha<\beta<\infty$ is treated in a similar way.) Then $\operatorname{fr}(I)=\{\beta\}$. We set $I_{1}:=(-\infty,\beta)$ and now use a well-known fact:
@@ 3. Bounds on the Derivatives of Bump Functions
By a bump function we mean here a $\mathcal{C}^{\infty}$-function $\alpha\colon{\mathbb{R}}\to{\mathbb{R}}$ which is 0 on $(-\infty,0]$, strictly increasing on $[0,1]$, and $1$ on $1,+\infty)$. Such bump functions play an important role in many constructions in analysis, e.g., partitions of unity. In this section we give a construction of a bump function $\alpha$ with controlled growth of its derivatives $\alpha^{(n)}$. As an auxiliary result we first establish the following formula for the derivatives of reciprocals of nonzero elements of a differential field $K$:
@@
We work in the setting of [4, Chapter 12] and assume familiarity with the concepts and the notations introduced there. Let $\phi$ range over $K^{\times}$ and iijjkkll over $\mathbb{N}$}. Put
@@
By the generalized Leibniz Rule, $\phi\mapsto h_{\phi}\colon K^{\times}\to K[[z]]^{\times}$ is a group morphism, so we obtain a group morphism $\phi\mapsto[h_{\phi}]\colon K^{\times}\to\mathfrak{tr}_{K}$ whose image is contained in the Appell group $\mathcal{A}$ over $K$ (cf. [4, p. 565]). For $i\leqslant j$ we have $h_{\phi}]_{ij}={j\choose i}\phi^{(j-i)}/\phi$. Since $\mathcal{A}\subseteq 1+\mathfrak{tr}_{K}^{1}$, we have $\big{(}[h_{\phi}]-1\big{)}^{k}\in\mathfrak{tr}_{K}^{k}$ for $k\geqslant 1$, and
$h_{\phi^{-1}}]=[h_{\phi}]^{-1}=\sum_{k}(-1)^{k}\big{(}[h_{\phi}]-1\big{)}^{k}=\sum_{k}(-1)^{k}\left(\sum_{l=0}^{k}{k\choose l}(-1)^{k-l}[h_{\phi^{l}}]\right).$
@@
Let $\alpha$, $c$, $d$ be as in Proposition 3.2, where we may assume $c\geqslant 1$. With $0^{0}:=1$, put $C_{n}:=cn^{dn}$, so $1=C_{0}\leqslant C_{1}\leqslant\cdots$. For $a<b$ in $\mathbb{R}$}, we define the increasing $\mathcal{C}^{\infty}$-function $\alpha_{a,b}\colon{\mathbb{R}}\to{\mathbb{R}}$ by
@@
Then $\alpha_{a,b,a_{*},b_{*}}(t)=0$ if $t\notin[a,b_{*}]$, $\alpha_{a,b,a_{*},b_{*}}(t)=1$ if $t\in[b,a_{*}]$, and $\alpha_{a,b,a_{*},b_{*}}$ is increasing on $[a,b]$ and decreasing on $a_{*},b_{*}]$. Suppose $\varepsilon,\varepsilon_{*}\leqslant 1$. Then by (3.6),
(3.8) $\lVert\alpha_{a,b,a_{*},b_{*}}\rVert_{n}\ \leqslant\ 3^{n}C_{n}\max\left\{\frac{1}{(b-a)^{n}},\frac{1}{(b_{*}-a_{*})^{n}}\right\}.$
In particular, for suitable reals $c_{1},d_{1}\geqslant 1$ (depending only on $bab_{*}-a_{*}$ccdd) we have $\lVert\alpha_{a,b,a_{*},b_{*}}\rVert_{n}\leqslant c_{1}n^{d_{1}n}$ for each $n$.
@@ 4. Weierstrass Transforms
Let $f\in\mathcal{C}^{m}({\mathbb{R}})$ be such that $\operatorname{supp}f$ (= the closure in $\mathbb{R}$} of the set ${\mathbb{R}}\setminus f^{-1}(0)$) is bounded; let also $\lambda$ range over $\mathbb{R}$$^{>}$. The function $f_{\lambda}\colon{\mathbb{R}}\to{\mathbb{R}}$} given by
@@
is known as the (generalized) Weierstrass transform with parameter $\lambda$; below we sometimes denote $f_{\lambda}$ also by $W_{\lambda}(f)$. We could have replaced the bounds -$\infty\infty$ in this integral by any aabb such that $\operatorname{supp}(f)\subseteq[a,b]$. A change of variables gives